OpenStudy (anonymous):

Show that the derivative of 1/tan(x) = -cosec^2 (x)

OpenStudy (loser66):

\[tan x =\frac{sin}{cos}\] so what is \(\dfrac{1}{tan x}\) =?

OpenStudy (anonymous):

\[\frac{ 1 }{\tan \theta}=-\csc ^{2} \theta \]

OpenStudy (anonymous):

cos / sin

OpenStudy (loser66):

sure, just take derivative of a fraction, apply quotient rule. that'sit

OpenStudy (anonymous):

thanks